OpenStudy (anonymous):

If c & d= c^2+ d^2-4, then 2 & 1= a.-2 b.-1 c.0 d.1 e.2

OpenStudy (anonymous):

If c & d= c^2+ d^2-4, then 2 & 1 that means c=2 and d=1 since c & d and 2 & 1 are in a same form. so just substitute 2 into the place of c and 1 into the place of d then it becomes: 2^2+1^2-4=4+2-4=2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[2 and 1=2^{2}+1^{2}-4=4+1-4=1\]

OpenStudy (anonymous):

oh yeah oppps mistake lol

OpenStudy (anonymous):

Sorry for that, made a calculation mistake D: